OpenStudy (anonymous):

-0.7x-0.16y=-4 5x+2y=20 solve for x and y

OpenStudy (triciaal):

one method is by substitution using the second eqn y = (20-5x) / 2 using this value of y in the first eqn -0.7x-0.16(20-5x)/2 = -4 -0.7x -1.6+0.4x = -4 solve for x with this value of x put in either of the original eqn and solve for y

OpenStudy (triciaal):

x=8 an y = -10 use any original eqn to verify the solution check in eqn 1 left side -0.7(8) -0.16(-10) = -5.6 --1,6 = -5.6 + 1.6 = -4 right side = -4 if you use eqn 2 5x + 2y = 20 5(8)-2(-10) =20 40-20 = 20 20 = 20 therefore it is verified and correct